OpenStudy (anonymous):

hi, how do i figure out 8 over 3 divided by negative 4?

OpenStudy (earthcitizen):

yh, it makes no sense to me. How can a positive number be divided by a - number, however \[8/3\div-4=8/3\times(-1/4)=-2/3\] this is what we all believe can you imagine that?